# Transition Path Sampling with Improved Off-Policy Training of Diffusion Path Samplers

### v4

@@ 3.3 Parameterization for high dimensional tasks
where $\bm{s}_{\theta}(\cdot)\in\mathbb{R}_{+}^{3N}$ is a neural network constrained to have positive output elements. One can observe the positive inner product between the bias force and the direction to minimize the Euclidean distance, i.e., $(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$, as follows:
$\bm{b}(\bm{X}_{t})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})=(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})^{T}\text{diag}(\bm{s}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t}))(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})\geq 0,$ (12)
which implies that the bias force can be decomposed into a vector positively aligning with $(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$ and a vector orthogonal to $(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$.
We also note that the bias force is $SE(3)$ equivariant to the current state $\bm{X}_{t}$ since it is the product of an invariant term $\text{diag}(\bm{s}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t}))$ and an equivariant term $(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$. This stems from the fact that the Kabsch algorithm is equivariant to the current state $\bm{X}_{t}$, i.e., roto-translating position $\bm{R}_{t}\mapsto\bar{\rho}\cdot\bm{R}_{t}$ leads to roto-translating output of the Kabsch algorithm $\rho_{t}\mapsto\bar{\rho}\cdot\rho_{t}$.
@@
In the experiments, we also consider other equivariant parameterizations that are less constrained: (1) directly predicting the bias force by $\rho_{t}\cdot\bm{b}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t})\in\mathbb{R}^{3N}$ and (2) predicting the force as gradient of the invariant bias potential $\nabla b_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t})\in\mathbb{R}$. We observe the parameterizations to be useful for low-dimensional tasks, but suffer from the sparse training signal on high-dimensional tasks.
@@
Indicator function parameterization. We also propose to relax the indicator function $1_{\mathcal{B}}$ as a radial basis function (RBF) kernel $\tilde{1}_{\mathcal{B}}(\bm{X})=k(\bm{R}_{T},\rho^{-1}\cdot{\bm{R}}_{\mathcal{B}};\sigma^{2})$ which measures the similarity between two positions and $\sigma>0$ controls the degree of relaxation. The range of RBF kernel $k$ is bounded by the interval $(0,1]$ so that $\log\tilde{1}_{\mathcal{B}}(\bm{X})$ is well-defined and $\tilde{1}_{\mathcal{B}}(\bm{X})$ represents the binary indicator function smoothly. To capture high training signal, we focus on the signal from subtrajectories of sampled paths, taking maximum over RBF kernel values of all intermediate states by $\tilde{1}_{\mathcal{B}}^{\text{max}}(\bm{X})=\max_{t\in[0,T]}k(\bm{R}_{t},\tilde{\bm{R}}_{\mathcal{B}};\sigma^{2})$. To extract the subtrajectory with a high training signal, we can truncate the paths at the time that maximizes RBF kernel values, allowing variable path lengths. Also, the relaxed indicator function is $SE(3)$ invariant to $\bm{R}_{t}$ due to the Kabsch algorithm.
@@ 4 Experiment
Table 1: Benchmark scores on the double-well system and three real-world molecular systems: Alanine Dipeptide, Polyproline Helix, and Chignolin. All metrics are averaged over 1024 paths for the double-well system, and 64 paths for real-world molecular systems. ETS is computed for paths that hit the target meta-stable state, and the best results are highlighted in bold. TPS-DPS predicting the bias force, potential, and atom-wise scaling are denoted by (F), (P), and (S), respectively. UMD ($\lambda$) denotes unbiased MD with temperature $\lambda$ and SMD ($k$) denotes steered MD with the force constant $k$. Unless otherwise specified, paths are generated by MD simulation at 1200K for double-well and 300K for real-world molecules.
@@
Polyproline Helix
UMD (300K) 2.22 $\pm$ 0.11 0.00 -
UMD (1200K) 1.38 $\pm$ 0.45 10.94 1010.28 $\pm$ 38.44
SMD (5k) 1.68 $\pm$ 0.17 54.69 350.58 $\pm$ 14.36
SMD (10k) 1.26 $\pm$ 0.06 100.00 355.62 $\pm$ 14.83
PIPS (F) 2.64 $\pm$ 0.15 0.00 -
PIPS (P) 1.85 $\pm$ 0.13 93.75 574.66 $\pm$ 20.49
TPS-DPS (F, Ours) 1.53 $\pm$ 0.12 98.44 418.17 $\pm$ 45.54
TPS-DPS (P, Ours) 1.35 $\pm$ 0.12 100.00 345.00 $\pm$ 32.58
TPS-DPS (S, Ours) 1.17 $\pm$ 0.02 100.00 342.00 $\pm$ 20.28
@@
TPS-DPS (S, Ours) 1.17 $\pm$ 0.66 59.38 -780.18 $\pm$ 216.93
In this section, we compare our method, called TPS-DPS, with both classical and ML approaches, assessing the accuracy and diversity of sampled transition paths. We begin with a synthetic double-well system at 1200K, followed by three real-world molecular systems with various numbers of amino acids: Alanine Dipeptide, Polyproline Helix, and Chignolin. Additionally, we conduct ablation studies to validate the effectiveness of each component in our method. All real-world molecular systems are simulated using the OpenMM library (Eastman et al., 2023). Details on OpenMM simulation and model configurations are provided in Sections B.1 and B.2, respectively.
Evaluation Metrics. We consider three metrics for evaluation: RMSD, THP, and ETS. The root mean square distance (RMSD) calculates the Euclidean distance between the atoms in the final position of the sampled path and the target position $\bm{R}_{\mathcal{B}}$, with the final position aligned to the target. The target hit percentage (THP) measures the number of paths that successfully arrive at the target meta-stable state $\mathcal{B}$. Finally, the energy of the transition state (ETS) measures the ability to identify probable transition states. For further details, refer to Section B.3.
Baselines. We compare TPS-DPS with both non-ML and ML baselines. For non-ML baselines, we consider unbiased MD (UMD) in various temperatures and steered MD (SMD; Izrailev et al., 1999) with various force constants $k$ and collective variables. For ML baselines, we consider a CV-free transition path sampling method, path integral path sampling (PIPS; Holdijk et al., 2024) which also trains a bias force by minimizing the KL divergence between path measures induced by the biased MD and the target path measure. For simplicity, we denote parameterizations for predicting force, potential, and atom-wise scaling as (F), (P) and (S), respectively.
(a) Double-well energy landscapes
(b) Distributions of the energy and $y$
Figure 2: Visualization of potential energy landscapes and distributions of the double-well system. (2(a)) Visualization of the neural bias potential $b_{\theta}$ for TPS-DPS (P) in the double-well potential energy landscape. (2(b)) Distributions of the potential energy (left) and $y$ coordinate (right) of transition states from 1024 transition paths sampled by each method.
4.1 Double-well system
(a) UMD
(b) SMD ($0.5$)
(c) SMD ($1$)
(d) TPS-DPS (F)
(e) TPS-DPS (P)
(f) Ground Truth
Figure 3: 16 transition path sampled from each method on the potential energy landscape of the double-well system. White circles indicate the meta-stable states and white stars indicate the saddle points. We sample paths from the left meta-stable state to the right. SMD ($k$) denotes steered MD with the force constant $k$. All paths are generated by MD simulation at 300K. SMD ($k$) denotes steered MD with the force constant $k$.
We begin by evaluating our method on a two-dimensional synthetic system, i.e., a double-well potential at 1200K. This system has two global minima (white circles) representing the meta-stable states, and two reaction pathways via saddle points (white stars). We sample transition paths from the left meta-stable state $\bm{R}_{\mathcal{A}}$ to the right meta-stable states =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\bm{R}-\bm{R}_{\mathcal{B}}\rVert<0.5\}$, using over-damped Langevin dynamics. We collect ground truth path ensembles by rejection sampling which proposes paths sampled from the unbiased MD simulations and accepts if the final states are in the target meta-stable states $\mathcal{B}$. We provide more details on the system in Section B.4.
In Table 1, TPS-DPS shows superior performance for the double-well potential regardless of the bias force design. In Figure 2, the neural bias potential accelerates the transition by increasing the potential energy near the initial meta-stable state while decreasing the potential energy near the two energy barriers. Moreover, the distribution of energy and $y$ coordinates of the transition states from TPS-DPS is closest to the ground truth compared with other baselines, successfully capturing two reaction channels. In Figure 3, unbiased MD at 1200K fails to escape the initial state while SMD struggles to pass the saddle points. Our approach generates transition paths more similar to the ground truth than other methods.
(a) UMD
(b) SMD (10)
(c) SMD (20)
(d) PIPS (F)
(e) TPS-DPS (F)
(f) TPS-DPS (P)
Figure 4: 16 sampled paths from each method on the Ramachandran plot of Alanine Dipeptide. White circles indicate meta-stable states and white stars indicate saddle points. We sample transition paths from the meta-stable state $C5$ (upper left) to $C7ax$ (lower right). SMD ($k$) denotes steered MD with the force constant $k$. Paths are generated by MD simulation at 1200K.
Figure 5: Snapshots of transition path for Alanine Dipeptide (left) and Polyproline Helix (right).
(a) UMD
(b) SMD (10k)
(c) SMD (15k)
(d) PIPS (P)
(e) TPS-DPS (P)
(f) TPS-DPS (S)
Figure 6: 16 sampled paths from each method projected to the top two TICA components. White circles indicate meta-stable states. We sample transition paths from the unfolded state (right) to the folded state (left). SMD ($k$) denotes steered MD with the force constant $k$. All paths are generated by MD simulation at 300K. (P) and (S) refer to predicting the bias potential and atom-wise scaling.
4.2 Alanine Dipeptide conformation change
We first consider Alanine Dipeptide for a real-world molecule consisting of two alanine residues, sampling transition paths from the $C5$ (upper left) to the $C7ax$ (lower right) as seen in Figure 4. The target meta-stable states are defined as =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\xi(\bm{R})-\xi(\bm{R}_{\mathcal{B}})\rVert<0.75\}$, where $\xi(\bm{R})=(\phi,\psi)$ is a well-known collective variable which consists of two backbone dihedral angles. Also, Alanine Dipeptide has two reaction channels between the $C5$ and $C7ax$ state passing through the saddle points marked as white stars.
In Table 1 and Figure 4, unbiased MD at 300K fails to escape the initial state, while SMD with the collective variable $\xi$ generates transition paths with less probable transition states. PIPS generates transition paths of only one reaction channel, suffering from mode collapse. Compared to the baselines, our method successfully generates diverse transition paths that pass two reaction channels and show superior performance regardless of the bias force design. For qualitative analysis, we plot the snapshots of the transition path sample from TPS-DPS in Figure 5.
4.3 Polyproline Helix isomerization
Next, we consider a more complex molecule, Polyproline Helix consisting of three proline residues. We sample transition paths from the left-handed state (PPII) in the cis-configuration to the right-handed state (PPI) in the trans-configuration. For simplicity, we define the target meta-stable state based on handedness $H$ as =$\mathcal{B}=\{\bm{R}~{}|~{}H(\bm{R})>0\}$. For the formal definition of handedness, refer to Section B.4. As seen in Table 1, our models consistently outperform all baselines: unbiased MD, SMD, and PIPS. In Figure 7(a), a transition path sampled from TPS-DPS crosses the energy barrier around 0.4ps and changes handedness for 1.5ps. Additionally in Figure 5, we plot the snapshots of the sampled transition path.
(a) Potential energy and CVs over time.
(b) Visualization of hydrogen bond formation in Chignolin.
Figure 7: Qualitative evaluation on transition path sampled from TPS-DPS. (7(a)) Potential energies and collective variables (CVs) of Alanine dipeptide, Polyporline Helix, and Chignolin from top to bottom. For CVs, we plot the distance of the backbone dihedral angles between the current state and the target state for Alanine Dipeptide, the handedness of the backbone for Polyproline Helix, and the hydrogen bond distance for Chignolin. One can see that the sampled path successfully transitioned to the target meta-stable state. (7(b)) Visualization of hydrogen bond formation for Chignolin in a sampled path. We project a sampled path on the top two TICA components. We visualize the time steps indicated by the black circles, with each hydrogen bond highlighted in green and yellow.
4.4 Chignolin folding

### v5

@@ 3.3 Parameterization for large systems
where $\bm{s}_{\theta}(\cdot)\in\mathbb{R}_{+}^{3N}$ is a neural network constrained to have positive output elements and predicts atom-wise scaling factors and the optimal roto-translation $\rho_{t}=\operatorname{argmin}_{\rho\in SE(3)}\lVert\bm{R}_{t}-\rho\cdot\bm{R}_{\mathcal{B}}\rVert$ which aligns $\bm{R}_{\mathcal{B}}$ with $\bm{R}_{t}$, as in Equation 2. We note that the bias force (divided by atom-wise masses) is positively correlated with the direction to target state, i.e., $(\bm{b}(\bm{X}_{t})/\bm{m})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})>0$.
To formalize the benefit of the positive correlation between the bias force and the direction to the target state, one can prove that there always exists a small enough step size $\Delta t$ that decreases the distance between the current state $\bm{R}_{t}$ and the aligned target state $\rho^{\prime}_{t+\Delta t}\cdot{\bm{R}}_{\mathcal{B}}$, i.e.,
$\lVert\rho^{\prime}_{t+\Delta t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t+\Delta t}^{\prime}\rVert<\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert,$ (12)
where $\bm{R}_{t+\Delta t}^{\prime}=\bm{R}_{t}+\bm{b}(\bm{X}_{t})\Delta t/\bm{m}$ is the position updated by the bias force with step size $\Delta t$ and $\rho^{\prime}_{t+\Delta t}=\operatorname{argmin}_{\rho\in SE(3)}\lVert\bm{R}^{\prime}_{t+\Delta t}-\rho\cdot\bm{R}_{\mathcal{B}}\rVert$. We provide the proof of Equation 12 in Section A.3.
@@
In the experiments, we also consider other equivariant parameterizations that are less constrained: (1) directly predicting the equivariant bias force $\bm{b}(\bm{X}_{t})=\rho_{t}\cdot\bm{b}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t})\in\mathbb{R}^{3N}$ and (2) predicting the invariant bias potential $b_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t})\in\mathbb{R}$ to obtain the bias force $\bm{b}(\bm{X}_{t})=-\nabla b_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t})\in\mathbb{R}^{3N}$. We observe these two parameterizations to be useful for low-dimensional tasks but struggle to produce meaningful paths in large systems during training. As shown in Figure 2, bias forces with the positive scaling parameterization are positively correlated with the direction to the target position regardless of network parameters, unlike direct force parameterizations.
(a) Direct prediction
(b) Positive scaling
Figure 2: Visualization of the bias force fields of two different bias force parameterizations with initialized neural networks. (2(a)) directly predicting the bias force and (2(b)) predicting the positive scaling factors of direction to the target position (white circle).
@@
Indicator function parameterization. We propose to relax the indicator function $1_{\mathcal{B}}$ into a radial basis function (RBF) kernel $\tilde{1}_{\mathcal{B}}(\bm{X})=k(\bm{R}_{T},\rho^{-1}\cdot{\bm{R}}_{\mathcal{B}};\sigma^{2})$ which measures the similarity between two positions where $\sigma>0$ controls the degree of relaxation. The range of RBF kernel $k$ is bounded by the interval $(0,1]$ so that $\log\tilde{1}_{\mathcal{B}}(\bm{X})$ is well-defined and $\tilde{1}_{\mathcal{B}}(\bm{X})$ represents the binary indicator function smoothly. To capture a high training signal from subtrajectories of sampled paths, we propose to take maximum over RBF kernel values of all intermediate states by $\tilde{1}_{\mathcal{B}}^{\text{max}}(\bm{X})=\max_{t\in[0,T]}k(\bm{R}_{t},\tilde{\bm{R}}_{\mathcal{B}};\sigma^{2})$. To extract the subtrajectory with a high training signal, we can truncate the paths at the time that maximizes RBF kernel values, allowing variable path lengths. Notably, the relaxed indicator function is $SE(3)$ invariant to $\bm{R}_{t}$ because of the Kabsch algorithm.
@@ 4 Experiment
Table 1: Benchmark scores on the double-well system and Alanine Dipeptide. All metrics are averaged over 1024 paths for the double-well system, and 64 paths for Alanine Dipeptide. ETS is computed for paths that hit the target meta-stable state, and the best results are highlighted in bold. Predicting the bias force, potential, and atom-wise positive scaling are denoted by (F), (P), and (S), respectively. UMD ($\lambda$) denotes unbiased MD with temperature $\lambda$ and SMD ($k$) denotes steered MD with the force constant $k$. Unless otherwise specified, paths are generated by MD simulation at 1200K for double-well and 300K for Alanine Dipeptide.
@@
In this section, we compare our method, termed TPS-DPS, with both classical non-ML and ML approaches, assessing the accuracy and diversity of sampled transition paths. We begin with the double-well system and Alanine Dipeptide followed by the four fast-folding proteins: Chignolin, Trp-cage, BBA, and BBL. Additionally, we conduct ablation studies to validate the effectiveness of each component in our method. All real-world molecular systems are simulated using the OpenMM library (Eastman et al., 2023). Details on OpenMM simulation and model configurations are provided in Sections B.1 and B.2, respectively. In Appendix C, we analyze the time complexity of TPS-DPS and evaluate the number of energy evaluations and runtime in training and inference time.
Evaluation Metrics. We consider three metrics to evaluate models: (Kabsch) RMSD, THP, and ETS. The root mean square distance (RMSD) measures the ability of the model to produce final positions of paths close to the target position $\bm{R}_{\mathcal{B}}$. The target hit percentage (THP) measures the ability of the model to produce final positions of paths that successfully arrive at the target meta-stable state $\mathcal{B}$. Finally, the energy of the transition state (ETS) measures the ability of the model to identify probable transition states. For further details, refer to Section B.3.
Baselines. We compare TPS-DPS with both non-ML and ML baselines. For non-ML baselines, we consider unbiased MD (UMD) with various temperatures and steered MD (SMD; Schlitter et al., 1994; Izrailev et al., 1999) with various force constants and collective variables (CVs). For ML baselines, we consider a CV-free transition path sampling method, path integral path sampling (PIPS; Holdijk et al., 2024) which also trains a bias force by minimizing the KL divergence between path measures induced by the biased MD and the target path measure.
(a) Potential energy landscapes
(b) Distributions of the energy and $y$
Figure 3: Visualization of potential energy landscapes and distributions in double-well potential. (3(a)) Visualization of the learned bias potential $b_{\theta}$ of TPS-DPS (P). (3(b)) Distributions of the potential energy and $y$ coordinate of transition states from 1024 transition paths sampled by each method.
4.1 Double-well system
UMD
SMD ($0.5$)
SMD ($1$)
TPS-DPS (F)
TPS-DPS (P)
Ground Truth
UMD
SMD (10)
SMD (20)
PIPS (F)
TPS-DPS (F)
TPS-DPS (P)
Figure 4: Visualization of sampled paths on energy landscapes. For the double-well system, we aim to sample transition paths from the left meta-stable state to the right on the potential energy landscape (top). For Alanine Dipeptide, we aim to sample conformational changes from the $C5$ (upper left) to the $C7ax$ (lower right) on the Ramachandran plot (bottom). White circles and stars indicate meta-stable states and saddle points, respectively.
We begin by evaluating our method on a synthetic two-dimensional double-well system with dual channels. This system has two global minima representing the meta-stable states and two reaction pathways with the corresponding saddle points. We aim to sample transition paths from the left state $\bm{R}_{\mathcal{A}}$ to the right meta-stable states =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\bm{R}-\bm{R}_{\mathcal{B}}\rVert<0.5\}$. We collect ground truth path ensembles by rejection sampling which proposes paths sampled from the unbiased MD simulations and accepts if the final states are in the target meta-stable states $\mathcal{B}$. We provide more details on the double-well system in Section B.4.
As shown in Table 1, Figure 3, and Figure 4, TPS-DPS outperforms baselines regardless of the bias force parameterizations and generates more similar transition paths to the ground truth than baselines. In Figure 3, the bias potential accelerates the transition by increasing the potential energy near the initial meta-stable state while decreasing the potential energy near the two energy barriers. Moreover, the distribution of energy and $y$ coordinate of the transition states from TPS-DPS is closest to the ground truth compared with other baselines, successfully capturing two reaction channels. In Figure 4, TPS-DPS (F) and (P) generate similar transition paths to the ground truth while UMD at 1200K fails to escape the initial state and SMD struggles to pass the saddle points.
4.2 Alanine Dipeptide
For real-world molecules, we first consider Alanine Dipeptide consisting of two alanine residues and aim to sample conformational changes from the meta-stable state $C5$ to $C7ax$ as shown in Figure 4. The target meta-stable states are defined as =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\xi(\bm{R})-\xi(\bm{R}_{\mathcal{B}})\rVert<0.75\}$, where $\xi(\bm{R})=(\phi,\psi)$ is a well-known collective variable which consists of two backbone dihedral angles. Alanine Dipeptide has two reaction channels with the corresponding saddle points.
In Table 1 and Figure 4, our method shows superior performance regardless of the bias force parameterizations and successfully generates diverse transition paths that capture two reaction channels. Compared to our method, UMD at 300K fails to escape the initial state, SMD with the two backbone torsion CV generates transition paths with less probable transition states, and two-way shooting struggles to find plausible transition states. PIPS generates transition paths with only one reaction channel, suffering from mode collapse.
Table 2: Benchmark scores on fast-folding proteins. All metrics are averaged over 64 paths. Unless otherwise specified, paths are generated at 300K for Chignolin and 400K for the others.
Method RMSD ($\downarrow$) THP ($\uparrow$) ETS ($\downarrow$)
Å % $\mathrm{kJmol}^{-1}$
@@ 4.2 Alanine Dipeptide
TPS-DPS (S, Ours) 1.17 $\pm$ 0.66 59.380 -780.18 $\pm$ 216.93
BBA
UMD (400K) 10.03 $\pm$ 0.390 0.00 -
UMD (1200K) 10.81 $\pm$ 1.050 0.00 -
SMD (10k) 2.89 $\pm$ 0.32 0.00 -
SMD (20k) 1.66 $\pm$ 0.30 26.560 -3104.95 $\pm$ 97.570
PIPS (F) 9.84 $\pm$ 0.18 0.00 -
PIPS (P) 9.09 $\pm$ 0.36 0.00 -
TPS-DPS (F, Ours) 9.48 $\pm$ 0.18 0.00 -
TPS-DPS (P, Ours) 3.89 $\pm$ 0.35 0.00 -
TPS-DPS (S, Ours) 1.21 $\pm$ 0.09 84.380 -3801.68 $\pm$ 139.38
Method RMSD ($\downarrow$) THP ($\uparrow$) ETS ($\downarrow$)
Å % $\mathrm{kJmol}^{-1}$
Trp-cage
UMD (400K) 7.94 $\pm$ 0.65 0.00 -
UMD (1200K) 8.27 $\pm$ 1.13 0.00 -
SMD (10k) 1.68 $\pm$ 0.23 3.12 -312.54 $\pm$ 20.67
SMD (20k) 1.20 $\pm$ 0.20 42.190 -226.40 $\pm$ 85.59
PIPS (F) 7.47 $\pm$ 0.19 0.00 -
PIPS (P) 6.07 $\pm$ 0.26 0.00 -
TPS-DPS (F, Ours) 6.35 $\pm$ 0.31 0.00 -
TPS-DPS (P, Ours) 3.15 $\pm$ 0.52 12.500 -512.97 $\pm$ 56.89
TPS-DPS (S, Ours) 0.76 $\pm$ 0.12 81.250 0-317.61 $\pm$ 140.89
BBL
UMD (400K) 18.48 $\pm$ 0.630 0.00 -
UMD (1200K) 18.90 $\pm$ 1.160 0.00 -
SMD (10k) 3.67 $\pm$ 0.22 0.00 -
SMD (20k) 2.97 $\pm$ 0.33 7.81 -1738.57 $\pm$ 386.81
PIPS (F) 17.92 $\pm$ 0.290 0.00 -
PIPS (P) 12.67 $\pm$ 0.310 0.00 -
TPS-DPS (F, Ours) 10.15 $\pm$ 0.540 0.00 -
TPS-DPS (P, Ours) 6.45 $\pm$ 0.26 0.00 -
TPS-DPS (S, Ours) 1.60 $\pm$ 0.19 43.750 -3616.32 $\pm$ 213.66
(a) Alanine Dipeptide
(b) Chignolin
(c) Visualization of Chignolin
Figure 5: Qualitative evaluation on transition path sampled from TPS-DPS. (5(a)) Potential energy and the two backbone dihedral angle distances between the current and target states. (5(b)) Potential energy and the two hydrogen bond distances between the current and the target state. (5(c)) Visualization of hydrogen bond formation in Chignolin. We highlight each hydrogen bond in green and yellow.
4.3 Fast-folding Proteins